OpenStudy (anonymous):

Steps to solve for 4x+8=x+2

OpenStudy (nottim):

make sure the constants and variables are on separate sides

OpenStudy (anonymous):

get the variables on one side and the constants on the other 4x+8 = x+2 -x = -x 3x+8 =2 -8 = -8 3x=-6 divide by 3 x=-2